40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Feb 24 17:42:00 2011, Aidan, Update, Green Locking, 15% of end laser sideband power transmitted through cavity 
    Reply  Sat Feb 26 02:33:16 2011, kiwamu, Update, Green Locking, sidebands on beatnote beat_note.png
       Reply  Sun Feb 27 09:43:59 2011, Aidan, Update, Green Locking, sidebands on beatnote 
       Reply  Sun Feb 27 13:09:56 2011, rana, Update, Green Locking, sidebands on beatnote 
Message ID: 4363     Entry time: Sun Feb 27 13:09:56 2011     In reply to: 4361
Author: rana 
Type: Update 
Category: Green Locking 
Subject: sidebands on beatnote 

When Koji and I were massaging the MC, we noticed that the oscillations were at 48.5 kHz. They were pretty huge and are probably what you're seeing on the beat. My guess is that they are the PZT resonances of the PSL 2W NPRO; we need to put a notch in the FSS box - it still has the notch from the old NPRO.

ELOG V3.1.3-